# JingFang: A Traditional Chinese Medicine Large Language Model of Expert-Level Medical Diagnosis and Syndrome Differentiation-Based Treatment

### v1

@@ Abstract
Traditional Chinese medicine (TCM) plays a vital role in health protection and disease treatment, but its practical application requires extensive medical knowledge and clinical experience. Existing TCM Large Language Models (LLMs) exhibit critical limitations of uncomprehensive medical consultation and diagnoses, and inaccurate syndrome differentiation-based treatment. To address these issues, this study establishes JingFang (JF): a novel TCM Large Language Model that demonstrates the expert-level capability of medical diagnosis and syndrome differentiation-based treatment. We innovate a Multi-agent Dynamic Collaborative Chain-of-Thought Mechanism (MDCCTM) for medical consultation, enabling JF with effective and accurate diagnostic ability. In addition, a Syndrome Agent and a Dual-Stage Retrieval Scheme (DSRS) are developed to significantly enhance the capacity of JF for disease treatment based on syndrome differentiation. JingFang not only facilitates the application of LLMs but also promotes the effective practice of TCM in human health protection and disease treatment.
Traditional Chinese Medicine, Large Language Model, Chain of Thought, LLM Agent, Retrieval-augmented Generation
@@
1 Introduction
Recently, the emergence of Large Language Models(LLMs) such as GPT-4(Achiam et al., 2023; Hurst et al., 2024), Llama3(Dubey et al., 2024), Deepseek-MOE(Dai et al., 2024a),Deepseek v2(Liu et al., 2024), and Qwen2.5(Yang et al., 2024a) has significantly advanced the field of natural language processing. These models not only possess exceptional language comprehension and generation capacity(Zhao et al., 2024), but also demonstrate strong cross-domain adaptability(Li et al., 2023; Bui et al., 2024; Thirunavukarasu et al., 2023) by integrating techniques such as pre-training(Han et al., 2021), fine-tuning(Ding et al., 2023), reinforcement learning(Cao et al., 2024), human feedback alignment(Yu et al., 2024), and knowledge graph retrieval(Wang et al., 2024). Among the numerous cross-domain applications of LLM, TCM has become one of the key focuses due to its importance in protecting human health and treating diseases(Lu et al., 2004; Zou et al., 2023; Ma et al., 2019), while the complexity of diagnosis and treatment of TCM poses a significant challenge.
In recent years, several studies have made some meaningful attempts, e.g. Zhongjing(Yang et al., 2024b), TCMChat(Dai et al., 2024b), Qibo (Zhang et al., 2024), BianCang (Wei et al., 2024), MedChatZH(Tan et al., 2024), etc. However, the critical limitations of the effective application of LLM in the diagnosis and treatment of TCM still remain. Unlike Western medicine, the core characteristic of TCM lies in its emphasis on treatment based on syndrome differentiation, which requires physicians to implement comprehensive medical diagnosis and accurately analyze symptoms and signs in patients. The diagnosis of TCM, which is the foundation for syndrome differentiation and treatment, is largely based on multiple rounds of personalized consultation to precisely collect patient information, which is complicated and uncertain. The scarcity of large-scale, accurately labeled multi-round consultation data leads to not good enough performance of existing models in extracting chief complaints of patients, mining symptoms, precise syndrome differentiation, and treatment. Additionally, current TCM LLMs lack explicit control over the diagnostic reasoning and decision process, which affects the model’s interpretability and reliability. Existing models still have limitations in adaptability, scalability, flexibility, and maintenance cost, making it difficult to quickly adapt to changing diagnostic scenarios or different TCM knowledge systems.
To address these critical challenges and limitations, the main contributions of this work are summarized below:
  1. 1.
    We develop JF: An expert-level medical diagnosis and syndrome differentiation-based treatment TCM LLM based on an innovative framework that integrates the technology of LLM Agent (Huang et al., 2024), Chain-of-Thought (CoT)(Chu et al., 2023), and Retrieval-Augmented Generation (RAG)(Gao et al., 2023) to significantly improve the completeness and precision of the model in medical consultation and syndrome differentiation-based treatment, addressing limitations of existing TCM models and empowering practical applications of LLM in TCM.
  2. 2.
    A Multi-agent Dynamic Collaborative Chain-of-Thought Mechanism (MDCCTM) is invented to enable JF with dynamic reasoning and explicit decision-making capabilities for medical consultation. According to the mechanism, multiple agents with various TCM specialties are established and dynamically collaborate to conduct professional consultation for comprehensive information collection and accurate syndrome differentiation-based treatment.
  3. 3.
    A Syndrome Agent and a Dual-Stage Retrieval Scheme (DSRS) is trained and developed correspondently based on the preconditioned multilevel knowledge of TCM, which significantly improves JF’s abilities of TCM syndrome differentiation and treatment in practical application
2 Framework of JingFang
In this section, we introduce the framework of the proposed JingFang (referred to as the JF framework). As shown in Figure 1, the JF framework is designed to consist of three main modules: TCM Consultation, TCM Syndrome Differentiation, and TCM Treatment Reconmmendation, according to the authentic process of TCM diagnosis and treatment. JF works well based on the process in the framework, following the designed MDCCTM and DSRS to fulfill the professional and accurate diagnosis and treatment.
TCM Consultation Module mainly involves the following steps:
① Expert Team Construction: The Medical Record Agent summarizes the patient’s symptoms and basic information for generating the chief complaint. Based on the chief complaint, the relevant TCM Specialist Agents and a TCM General Agent are gathered to form the TCM expert team for the diagnosis.
TCM Treatment Recommendation Module provides professional and targeted treatment recommendations for patients. The Treatment Agent is generated through a DRS to provide personalized treatment suggestions based on the patient’s syndrome type and the patient’s condition. For a detailed description of each module construction see Sec.3.
@@ 3 Method
In this section, we introduce the innovative and well-structured MDCCTM, DSRS, and the trained Syndrome Agent in detail.
3.1 Multi-agent Dynamic Collaborative Chain-of-Thought Mechanism
After an in-depth study of the TCM consultation process, we found that experienced TCM experts typically pre-set one or two follow-up consultation questions based on the patient’s known condition, and make flexible and targeted adjustments to them based on more information during the consultation process. Inspired by this, the MDCCTM is developed to better fit the actual TCM consultation process, and the prompts: $prompt_{[\ ]}$ here, are particularly designed to facilitate agents to implement the specific functions. The MDCCTM can be formulated as follows:
@@ 3.1.2 Consultation CoT Construction
To start the CoT construction at the t-$th$ round of consultation, the TCM Medical Record Agent summarizes the patient’s current condition $Con_{t}$ from the current multi-round consultation result that is defined as: $C_{t}=\{(Q_{k},A_{k})\mid k\in[0,t]\}$, where $Q_{k}$,$A_{k}$ is the question and the corresponding answer at k-$th$ round of consultation. Thus, $Con_{t}$ can be defined as:
@@
$ga_{t}=CoT_{gen}(C_{t},\ GD,\ prompt_{ga})$ (5)
The expert team’s consultation CoT is constructed by combining the $qa_{it}$ and $ga_{t}$ at t-$th$ consultation :
$TCoT_{t}=(qa_{ti},\ ga_{t})$ (6)
This consultation CoT lays the foundation for subsequent consultation processes. In the following steps, the expert team iteratively revises and optimizes $TCoT_{t}$ to make the consultation process more refined and targeted.
@@ 3.1.3 Consultation CoT Integration and Evaluation
Integrating the widely used classical framework Ten Questions Song(Hong et al., 2011) (TQS) that helps doctors assess the patient’s condition from multiple dimensions, we establish a Consultation CoT Evaluation Algorithm (CCEA).
@@
To evaluate the comprehensiveness of the questions, we calculate the similarity of questions in the consultation CoT with those in the TQS as the score of comprehensiveness by using an embedding model. To evaluate the pertinence of the questions, the key issues of each expert (different TCM experts focus on the characteristic diseases and diagnostic points related to their specialty, e.g., internal medicine focuses on organ function) and the patient’s known condition are incorporated as the basis for the pertinence score. The final score is obtained by aggregating these two scores to evaluate the consultation quality.
Based on the patient’s current condition $Con_{t}$ and $TCoT_{t}$, an Evaluation Agent is designed to summarize the consultation CoT provided by the expert team and evaluate the consultation questions within the summarized consultation CoT via the TCM knowledge-based CCEA. Thus, a holistic consultation CoT $RCoT_{t0}$ serves as the summary of the expert team’s consultation CoT that includes the score for each question generated as:
@@ 3.1.4 Consultation CoT Analysis and Optimization
Agents in the expert team further analyze the summary $RCoT_{t0}$ based on the evaluation results, and provide feedback on whether they agree with it or not. If an expert disagrees, it will provide modification suggestions for low-score questions and optimize the consultation CoT to get a higher score calculated by CCEA. Through this optimization and feedback process, the TCM expert team ensures that the consultation CoT is both comprehensive and targeted. Specifically, in $j$-th round of feedback, expert suggestions are recorded as $Mod_{j}$ and the summary $RCoT_{tj}$ is updated as follows:
@@
The integration and optimization process continues until all experts agree on the summary or the maximum number of feedback rounds is reached, a final Consultation CoT $RCoT_{tf}$ is generated to start the official consultation in this round.
@@ 3.1.5 Multi-Round Consultation
Based on the consultation CoT $RCoT_{tf}$ and the current multi-round consultation result $C_{t}$ in the t-$th$ round, a Consultation Agent is developed according to real-world consultation scenarios to generate the next round of consultation questions:
$Q_{t+1}=Q_{gen}(RCoT_{tf},\ Con_{t},\ prompt_{q}),$ (9)
and interact with the patient. After this, the Medical Record Agent is called to systematically integrate the patient’s condition into $Rec_{t}$, which is
@@
Following this process, a series of professional and targeted questions is generated to further explore the patient’s condition. When the key information collected in the record of the patient is completed or the maximum number of consultation rounds is reached according to the Consultation Information Collection Algorithm3, the Medical Record Agent presents the final case $Rec_{f}$ to the next Syndrome Differentiation and Treatment Recommendation Module.
Figure 2: An example of MDCCTM in the medical consultation
Fig.2 exhibits an example of how this MDCCTM works in a medical consultation scenario. For more detailed and specific application examples of this MDCCTM in the real medical consultation scenario see A in the appendix.
@@ 3.2 TCM Syndrome Agent
In the field of TCM, the accuracy of syndrome differentiation is the cornerstone of precise treatment. High-quality datasets are essential for LLMs to develop accurate syndrome differentiation capabilities. We observed that real-world TCM medical records and diagnostic data often contain a large amount of information unrelated to the patient’s condition, i.e., noise, which cannot be easily removed through conventional regularization techniques. For example, phrases such as “The patient visited our hospital for further integrated traditional Chinese and Western medicine treatment” or “The patient was advised to undergo cranial CT and blood tests but refused” are not only irrelevant to the core condition but also interfere with the model’s ability to focus on key disease-related information. At the same time, content in the medical record, such as “Blood amylase test showed no significant abnormalities and abdominal CT indicated slight thickening of the ascending colon wall”, which requires medical instrument examinations, also needs to be screened out.
According to the TQS, it helps TCM experts to comprehensively understand the patient’s symptoms, medical history, and physical signs through a series of fundamental questions, providing key information for precise diagnosis. Inspired by this, we propose a universal data preprocessing method by utilizing LLM to automatically extract key information related to the TQS from raw data and strictly retain core information closely related to the patient’s condition. This ensures alignment between the case outputs in the JF and the format of the training data cases, aiming to significantly enhance the accuracy of syndrome differentiation through more systematic and precise data processing.
To enable JF with TCM syndrome differentiation ability and smoothly complete accurate syndrome differentiation tasks within its framework, we particularly develop a TCM Syndrome Agent based on the fine-tuned LLM. Moreover, according to the final patient’s condition summary from the multi-round consultation $Rec_{f}$, the TCM Syndrome Agent is used to identify the most likely syndrome type for the patient:
@@
The detailed description of syndrome differentiation database preconditioning and model fine-tuning see Sec.B.1 and Sec.B.2 respectively in the appendix.
@@ 3.3 TCM Treatment Agent
In TCM diagnosis and treatment, treatment recommendations are essentially formulated from the perspective of syndrome differentiation. In addition, TCM experts also integrate various aspects of a patient’s information, including chief complaints, medical history, and constitution, dynamically adjusting treatment plans in response to changes in the patient’s condition to ensure precise treatment. This complicated and changeable procedure is a characteristic of TCM treatment but poses a challenge to the precise treatment of LLMs. To facilitate LLMs with accurate and targeted treatment recommendations ability, a Dual-Stage Retrieval Scheme (DSRS) is proposed to offer the exact and targeted treatment recommendation. Here is the algorithm of DSRS, and the detailed introduction of it see Sec.C in the appendix.
Algorithm 2 Dual-Stage Retrieval Scheme
1:  Initialization:$max\_results\leftarrow 3$
[…]
7:  Output: $hybrid\_results$
According to DSRS, a TCM Treatment Agent that utilizes the patient’s condition summaries and syndrome differentiation results derived from multiple rounds of consultation is developed to automatically screen and recommend the most suitable prescriptions from a structured TCM prescription database:
$TCM_{t}=Treat_{gen}(Rec_{f},DB_{TCM},prompt_{p}).$ (12)
$DB_{TCM}$ represents the delicately constructed TCM prescription database containing various TCM fields, including internal medicine, gynecology, pediatrics, and surgery. Each entry in the database includes essential information such as disease categories, syndrome types, clinical manifestations, representative formulas, commonly used herbs, and other therapeutic methods. Fig.3 illustrates an example of the performance of the TCM Treatment Agent, and for more information see Sec.E.4 in the appendix.
Figure 3: An example of the TCM Treatment Agent performance

### v2

# JingFang: An Expert-Level Large Language Model for Traditional Chinese Medicine Clinical Consultation and Syndrome Differentiation-Based Treatment

@@ Abstract
The effective application of traditional Chinese medicine (TCM) requires extensive knowledge of TCM and clinical experience. The emergence of Large Language Models (LLMs) provides a solution to this, while existing LLMs for TCM exhibit critical limitations of incomplete clinical consultation and diagnoses, as well as inaccurate syndrome differentiation. To address these issues, we establish JingFang (JF), a novel TCM LLM that demonstrates the level of expertise in clinical consultation and syndrome differentiation. We propose a Multi-Agent Collaborative Chain-of-Thought Mechanism (MACCTM) for comprehensive and targeted clinical consultation, enabling JF with effective and accurate diagnostic ability. In addition, a Syndrome Agent and a Dual-Stage Recovery Scheme (DSRS) are developed to accurately enhance the differentiation of the syndrome and the subsequent corresponding treatment. JingFang not only facilitates the application of LLMs but also promotes the effective application of TCM for healthcare.
1 Introduction
In recent years, the emergence of Large Language Models(LLMs) such as GPT-4o[1, 18], Llama3 [12], Deepseek-MOE [7],Deepseek v3 [10], Gemini-1.5-pro [23] and Qwen2.5 [30] has significantly advanced the field of natural language processing. These models not only possess exceptional language comprehension and generation capacity [33], but also demonstrate strong cross-domain adaptability [19, 3, 25] by integrating techniques such as pre-training [15], fine-tuning [11], reinforcement learning [4], human feedback alignment [32], and knowledge graph retrieval [27]. Among the numerous cross-domain applications of LLM, TCM has become one of the key focuses due to its importance for health care and disease treatment [20, 34, 21].
There are several studies have made some meaningful attempts, e.g., HuaTuo [26](a TCM diagnostic system based on LLM that focuses on simulating the syndrome differentiation thinking process of TCM physicians), TCMChat [8](the generative LLM for TCM built on Baichuan2-7B-Chat and trained on 6 task categorie diverse TCM data), Zhongjing [31] (the reinforcement learning with human feedback combines with CMtMedQA boosted the model’s proactive reasoning and knowledge alignment in complex TCM diagnostics), HuaTuo [5](It proposes an innovative one-stage domain adaptation protocol that unifies heterogeneous data into instruction-output pair format and introduces a data priority sampling strategy), ShenNong [28](It proposes a framework that improves large language models’ performance on TCM tasks using limited data through a two-stage approach: supervised fine-tuning with medical cases followed by reinforcement learning from AI feedback), and SunSimiao [29](the 7B model is fine-tuned from Qwen2-7B with high-quality medical data, achieving excellent performance equivalent to 30B-scale models on CMB-Exam).
However, the critical limitations of the effective application of LLM in TCM still remain. The core characteristic of TCM is the emphasis on syndrome differentiation and treatment, which requires physicians to conduct a comprehensive clinical consultation, accurately analyze patients’ symptoms and signs, and then carry out the corresponding treatment. Comprehensive clinical consultation is the foundation for syndrome differentiation and treatment, which largely relies on multiple rounds of professional and targeted inquiries to accurately collect information. The scarcity of large-scale and accurately labeled multi-round consultation data leads to not good enough performance of existing models in clinical consultation and diagnosis. They also lack explicit control over the clinical consultation process and diagnostic reasoning, affecting the interpretability and reliability of the models. In addition, existing models have significant limitations in the precise differentiation of TCM syndromes, which affects their effectiveness in disease treatment and their application in real clinical scenarios.
To address these, the main contributions of this work are summarized as:
  1. 1.
    We develop JF, a TCM LLM that addresses the limitations of existing models and empowers effective applications of LLM in TCM. JF’s innovative framework endows it with expert-level clinical consultation and syndrome differentiation capabilities.
  2. 2.
    The Multi-Agent Collaborative Chain-of-Thought Mechanism (MACCTM) is designed to enhance model’s comprehensive clinical consultation abilities, enabling professional and targeted clinical interactions while strengthening its explicit reasoning, decision and interpretability during the consultation process.
  3. 3.
    The Syndrome Agent has been systematically trained based on preprocessed and structured TCM knowledge, combining with the developed Dual-Stage Retrieval Scheme (DSRS) to substantially improve the model’s capabilities in precise syndrome differentiation and the corresponding treatment.
@@ 2 Method
2.1 Framework of JingFang
As shown in Figure 1, integrating the technology of LLM Agent  [17], Chain-of-Thought (CoT) [6], and Retrieval-Augmented Generation (RAG) [14], the framework of JF consists of three main modules: TCM Consultation, TCM Syndrome Differentiation, and TCM Treatment Reconmmendation, following the authentic process of TCM diagnosis and treatment. The MACCTM in TCM Consultation enables JF’s ability to provide comprehensive and targeted clinical consultation, the Syndrome Agent in TCM Syndrome Differentiation improves JF’s ability to accurately differentiate syndrome, and the DSRS in TCM Treatment Reconmmendation ensures the model to provide accurate treatment recommendations according to diagnosis and syndrome differentiation. See A in details.
2.2 Multi-Agent Collaborative Chain-of-Thought Mechanism
After an in-depth study of the TCM consultation process, we found that experienced TCM experts typically pre-set one or two follow-up questions based on the patient’s known situations, and make adaptive and targeted inquiries to collect more information during the consultation process. Inspired by this, the MACCTM is developed to better fit the authentic TCM consultation process. The prompts: $prompt_{[\ ]}$ here, are particularly designed to facilitate agents to implement a specific function, so the MACCTM can be formulated as follows:
@@ 2.2.2 Consultation CoT Construction
To start the CoT construction at the t-$th$ round of consultation, the TCM Medical Record Agent summarizes the patient’s current condition $Con_{t}$ from the current consultation result that is defined as: $C_{t}=\{(Q_{k},A_{k})\mid k\in[0,t]\}$, where $Q_{k}$ and $A_{k}$ is the question and the corresponding answer at k-$th$ round of consultation. Thus, $Con_{t}$ can be defined as:
@@
$ga_{it}=CoT_{gen}(C_{t},\ GD,\ prompt_{ga})$ (5)
The expert team’s consultation CoT is constructed by combining $qa_{it}$ and $ga_{it}$ at t-$th$ consultation:
$TCoT_{t}=(qa_{it},\ ga_{it})$ (6)
This consultation CoT lays the foundation for subsequent consultation process. In the following steps, the expert team iteratively revises and optimizes $TCoT_{t}$ to make the consultation process more professional and targeted.
@@ 2.2.3 Consultation CoT Integration and Evaluation
Integrating Ten Questions Song [16] (TQS), the widely used classical framework in TCM consultation, we establish a Consultation CoT Evaluation Algorithm (CCEA).
@@
We calculate the similarity of questions in the consultation CoT with those in the TQS as the score of comprehensiveness via an embedding model. For the evaluation the pertinence of the questions, the key issues of each expert (different TCM experts focus on the characteristic diseases and diagnostic points related to their specialty) and the patient’s known condition are incorporated to calculate their similarity as the basis for the pertinence score. The final score is obtained by aggregating the comprehensiveness score and the pertinence score as the evaluation the consultation. Thus, based on the patient’s current condition $Con_{t}$, an Evaluation Agent is designed to conclude the expert team’s consultation CoT $TCoT_{t}$ and evaluate the consultation questions via the TCM knowledge-based CCEA. Thus, a holistic consultation CoT $RCoT_{t0}$ that includes evaluation and summary for each question is generated:
@@ 2.2.4 Consultation CoT Analysis and Optimization
Agents in the expert team further analyze the $RCoT_{t0}$ based on the evaluation scores, and provide feedback on whether they agree with it or not. If an expert disagrees, it will provide modification suggestions for low-score questions and optimize the consultation CoT to get a higher score. Through this optimization and feedback process, the TCM expert team ensures that the consultation CoT is both professional and targeted. Specifically, in $j$-th round of feedback, expert suggestions are recorded as $Mod_{j}$ and the $RCoT_{tj}$ is updated as follows:
@@
The modification and optimization process continues until all experts agree on the summary or the maximum number of feedback rounds is reached, a final Consultation CoT $RCoT_{tf}$ is generated to start the formal consultation in this round.
@@ 2.2.5 Multi-Round Consultation
Based on the final consultation CoT $RCoT_{tf}$ and the current consultation result $C_{t}$ in the t-$th$ round, a Consultation Agent is developed following the real-world consultation scenarios to generate the next round of consultation question and interact with the patient:
$Q_{t+1}=Q_{gen}(RCoT_{tf},\ Con_{t},\ prompt_{q}).$ (9)
After this, the Medical Record Agent is called to systematically integrate the patient’s condition into medical record $Rec_{t}$, which is
@@
As the process continues, a series of professional and targeted questions is generated to further explore the patient’s condition. When the key information collected in the medical record is completed or the maximum number of consultation rounds is reached according to the Consultation Information Collection Algorithm2, the Medical Record Agent presents the final medical case $Rec_{f}$ to the Syndrome Differentiation Module and Treatment Recommendation Module.
Figure 2: An example of MACCTM in the medical consultation
Fig.2 exhibits an example of how this MACCTM works in a medical consultation scenario. For more detailed and specific examples of the MACCTM in the real medical consultation scenario, see B in the appendix.
@@ 2.3 TCM Syndrome Agent
In TCM, the accuracy of syndrome differention is fundamental for precise treatment that requires the high-quality datasets for LLM training. We found that real-world TCM records contain substantial noise resistant to conventional regularization. This noise includes irrelevant administrative statements, such as "patient visited for integrated treatment", and western medical examination references, such as "patient refused cranial CT", which distract models from essential disease information. Additionally, instrument-based diagnostic content, e.g., "blood amylase tests normal, abdominal CT shows colon wall thickening" must be filtered out to maintain focus on TCM-relevant symptom patterns.
In real TCM scenarios, TCM experts usually utilize TQS to comprehensively understand the patient’s symptoms, medical history, and physical signs through a series of fundamental questions, obtaining key information for precise diagnosis. Inspired by this, we propose a universal data preprocessing method by utilizing LLM to automatically extract key information related to the TQS from raw data and strictly retain core information closely related to the patient’s condition. This ensures alignment between the case outputs in the JF and the format of the training medical cases data. The systematic preprocessing of data has laid a solid foundation for subsequent accurate diagnosis and precise syndrome differentiation, significantly enhancing the model’s ability of syndrome differentiation-based treatment.
To enable JF with ability of precise syndrome differentiation, we particularly train and develop a TCM Syndrome Agent via fine tuning the LLM. Moreover, according to the final patient’s medical record from the multi-round consultation $Rec_{f}$, the TCM Syndrome Agent is used to identify the most likely syndrome type for the patient:
@@
For the detailed description of syndrome differentiation database preconditioning and the model fine tuning, see Sec.C.1 and Sec.C.2 in the appendix, respectively.
@@ 2.4 TCM Treatment Agent
In TCM diagnosis and treatment, treatment recommendations are essentially formulated from the perspective of syndrome differentiation. In addition, TCM experts also integrate various aspects of patient’s information, including chief complaints and medical history, dynamically adjusting treatment plans to ensure the precise and targeted treatment. This complicated and changeable procedure is a characteristic of TCM treatment but poses a challenge to the precise treatment of LLMs. To facilitate LLMs with accurate and targeted treatment recommendations ability, a Dual-Stage Retrieval Scheme (DSRS) is proposed. For more detailed information of this method, please refer to section D in the appendix.
According to DSRS, a TCM Treatment Agent that utilizes the patient’s medical record and syndrome differentiation result is developed to automatically screen and recommend the most suitable prescriptions from a structured TCM prescription database. It can be formulated as:
$TCM_{t}=Treat_{gen}(Rec_{f},DB_{TCM},prompt_{p}),$ (12)
where $DB_{TCM}$ represents the delicately constructed TCM prescription database containing various TCM fields, including internal medicine, gynecology, pediatrics, and surgery. Each entry in the database includes essential information such as disease categories, syndrome types, clinical manifestations, representative formulas, commonly used herbs, and other therapeutic methods. More information can be found in section F.4 in the appendix.